OpenStudy (january123):

Look at the relationship between a and b. a 3 7 11 15 b 27 63 99 135 Which equation below describes the relationship between a and b? b = 24 + a b = 81 ÷ a b = 9 × a b = 30 - a

OpenStudy (january123):

@pooja195 @apoorvk @alesia/bailey

OpenStudy (alesia/bailey):

b=9 X a 3*9=27 7*9=63 11*9=99 15**9=135

OpenStudy (january123):

THX

OpenStudy (alesia/bailey):

no problem ^^